OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality -2x - y >= -8 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

Can you solve the inequality? After that, keep in mind that: Less than(< or <=)... Shading to the LEFT Greater than(> or >=)... Shading to the RIGHT Equal sign(>= or <=)... CLOSED circle NO equal sign(> or <)... OPEN circle

OpenStudy (anonymous):

-2x - y >= -8 Add 2x -y >= -2x - 8 Divide by -1. FLIP SIGN y <= 2x + 8.

OpenStudy (anonymous):

closed would be solid and open would be dashed right??

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

ok now early i thought it was D

OpenStudy (anonymous):

am i right